OpenStudy (anonymous):

Find the solution to the system of equations. x+y+z=8 y+z=5 z=2

OpenStudy (anonymous):

A. x = 3, y = 3, z = 1 B. x = 3, y = 3, z = 2 C. x = 8, y = 3, z = 2 D. x = 4, y = 3, z = 2

OpenStudy (solomonzelman):

You are given the z, so substitute that into the other two equations.

OpenStudy (solomonzelman):

x + y + 2 = 8 y + 2 = 5

OpenStudy (anonymous):

It would be the A

OpenStudy (solomonzelman):

Now, the solution for y comes in handy from the 2nd equation, and then substitute the solution for y into the 1st equation and solve for x.

OpenStudy (solomonzelman):

It is impossible to be A, because you already know that z=2 (not 1).

OpenStudy (anonymous):

so it would be B @SolomonZelman ?

OpenStudy (solomonzelman):

Yes, B is correct !

OpenStudy (anonymous):

Ops wasn't paying attention I meant to say B thanks for the catch :)

OpenStudy (anonymous):

Thanks !